OpenStudy (anonymous):

Write the Expression as a sum and/or difference of logarithms. Express powers as factors. ln[(x^2-x-2)/(x+4)]^(1/3) x>2

OpenStudy (gorv):

ln[(x^2-x-2)/(x+4)]^(1/3) =(1/3)*[ln(x^2-x-2)-ln(x+4)]

OpenStudy (gorv):

log(a/b)=log a - log b

OpenStudy (anonymous):

Must it be simplified further or is that the final form?

OpenStudy (gorv):

final form it is